# On GRPO Collapse in Search-R1: The Lazy Likelihood-Displacement Death Spiral

### v1

@@ 5.1 Likelihood Dynamic
To illustrate the prevalence and progression of LLD during training, we visualize the likelihood-displacement trajectory in tool-integrated RL. As shown in Fig.˜2, the evolution unfolds in three characteristic phases. Phase I (early stagnation): the likelihood of correct responses remains nearly unchanged even as the reward increases, showing the initial emergence of LLD.
The later two phases correspond to the $\epsilon\leq 0$ regime, where the likelihood of correct responses strictly decays rather than improves. Phase II (steady-decay, steps 60–120): the likelihood exhibits a slow but consistent downward drift, while the reward grows only gradually and the gradient norm stays stable—indicating persistent LLD. Phase III (acceleration, after step 120): the likelihood begins to collapse sharply, coinciding with a rapid surge in gradient magnitude (marked by the red star), which triggers gradient explosion and ultimately leads to training collapse.
The zoomed-in view on the right highlights this transition: likelihood keeps decreasing while gradients accelerate upward, and the reward curve begins to fall. Although the final collapse is caused by exploding gradients, the likelihood decay is present throughout training, introducing an accumulating instability that becomes dramatically amplified during the acceleration phase (see Appendix Sec.˜B.1 for more discussion).
5.2 Lazy Likelihood Displacement Death Spiral
In Fig.˜2, we observe an accelerated decline in response likelihood as training progresses. In this section, we first simplify notation by writing the feedback-masked trajectory as $\hat{\bm{y}}=(\bm{y}_{0},\bm{y}_{1},\ldots,\bm{y}_{T})$. We then formally characterize this accelerated likelihood decay as a Lazy Likelihood-Displacement (LLD) death spiral.
Definition 5.1 (LLD Death Spiral)
@@ 5.2 Lazy Likelihood Displacement Death Spiral
Consider a tuple $(\bm{x},\hat{\bm{y}}^{+})$ and a policy update from $\pi_{\theta_{t}}$ to $\pi_{\theta_{t+1}}$. We say the system enters an LLD Death Spiral if the policy evolution exhibits the following self-reinforcing progression:
$\mathrm{LLD}_{t}\;\Longrightarrow\;C^{\mathrm{low}}_{t}\;\Longrightarrow\;\mathrm{LLD}_{t+1},\quad\epsilon_{t+1}<\epsilon_{t}\leq 0$
where $\mathrm{LLD}_{t}$ denotes lazy likelihood displacement at time $t$ with decreasing likelihood ($\epsilon_{t}\leq 0$), and $C^{\mathrm{low}}_{t}$ denotes low-confidence trajectories whose likelihood decreases under $\pi_{\theta_{t}}$. The transition $\mathrm{LLD}_{t}\Longrightarrow\ C^{\mathrm{low}}_{t}$ reflects that degraded likelihood leads to increasingly diffuse predictions, while the transition $C^{\mathrm{low}}_{t}\Longrightarrow\ \mathrm{LLD}_{t+1}$ occurs because low-confidence incorrect responses often contain actions similar to those in correct responses, inducing a larger value in Eq.˜9 and thus a stronger negative-gradient influence. When these effects compound across iterations, the likelihood decay accelerates, creating a self-perpetuating collapse that we refer to as the LLD Death Spiral.
@@
Then we demonstrate the LLD Death Spiral through accelerated entropy explosion and accelerated per-sample LLD.
Figure 3: Effect of likelihood displacement across different training iterations for the Qwen2.5-3B-Instruct model. Results are computed on the first 50 samples of the training set, discarding cases where all responses are uniformly correct or uniformly incorrect. Bars below zero (orange) indicate samples whose correct responses’ likelihood decreases after training.
@@
Accelerated Per-Sample LLD. We conduct a controlled experiment on the searching-based question answering NQ (kwiatkowski2019natural) dataset to examine how the negative gradient of GRPO affects the log-likelihood of correct responses. Using Qwen2.5-3B-Ins (yang2024qwen2), we generate 8 rollouts per question and retain only those examples that contain a mixture of correct and incorrect responses, discarding cases where all responses are uniformly correct or incorrect. To isolate the per-sample learning dynamics, we reinitialize the model parameters $\theta$ for each individual question, apply a single GRPO update to obtain $\theta^{\prime}$, and then measure the average log-likelihood change of the correct responses:
$\Delta(\bm{x}):=\frac{1}{N^{+}}\sum_{i=1}^{N^{+}}\left[\ln\pi_{\theta^{\prime}}(\hat{\bm{y}}_{i}^{+}\mid\bm{x})-\ln\pi_{\theta}(\hat{\bm{y}}_{i}^{+}\mid\bm{x})\right],$ (3)
for a question $\bm{x}$ with $N^{+}$ correct responses. We present the results in Fig.˜3. In the early stage of training (iteration 50), only mild LLD is observed: most questions show negligible reductions in the likelihood of correct responses, likely because correct and incorrect trajectories remain structurally dissimilar. However, as training progresses, the effect becomes more pronounced. Increasing structural similarity between incorrect and correct responses leads to broader likelihood degradation, as shown by the rising orange curve. As training enters the acceleration phase (iterations 120–140), the trend intensifies into an LD death spiral, with over half of the samples at iteration 140 showing substantial drops in the likelihood of correct responses. This collapse is driven by the extremely low likelihood of incorrect responses, which amplifies negative-gradient contributions and destabilizes learning.
@@
Accelerated Entropy Explosion. Fig.˜4 visualizes the evolution of entropy, response length, and valid-search ratio during training for Qwen2.5-3B-Ins (a) and Qwen2.5-3B-Base (b). Across both models, mean token entropy initially increases slowly, mirroring the early slow-LLD regime in Fig.˜2, but later transitions into a phase of sharply accelerated growth that aligns with the acceleration phase of the LLD trajectory. This steepening entropy slope is a clear demonstration of the death spiral: as low-confidence responses accumulate, the model assigns increasingly diffuse, low-probability distributions to its tokens, which further strengthens LD and drives the system toward instability. Crucially, this entropy acceleration occurs while response length and valid-search counts remain nearly unchanged, confirming that the rising entropy and the corresponding likelihood decay is not caused by trajectory length or tool-query frequency, but is instead a direct manifestation of the LD effect itself.
Figure 4: We illustrate how entropy, response length, and valid-search ratio evolve during training. For both Qwen2.5-3B-Instruct (a) and Qwen2.5-3B-Base (b), entropy exhibits a accerlerrated upward trend prior to collapse indicating a strong LD issue. Meanwhile, the response length and valid-search times remain stable in the early stages but later begin to fluctuate markedly and eventually drop sharply.
5.3 Watch for correct actions embedded within incorrect responses.
Observing tool-integrated GRPO is substantially more susceptible to LLD than non-tool settings (deng2025effect). This section examines LLD at the granularity of response actions, where each action is defined by the segments between consecutive tool calls. Our analysis on Qwen2.5-3B-Ins (yang2024qwen2) reveals a striking pattern unique to tool-integrated GRPO: correct actions frequently appear within otherwise incorrect responses, and this intermixed structure disrupts the likelihood estimation process, amplifying LLD and destabilizing training.
Frequent correct actions in Incorrect responses. In our experiments, we found that the model typically uses its first response action to generate a search query, and the correctness of this initial action increases steadily as training progresses. To be specific, we measure whether the retrieved documents produced by an incorrect response match those retrieved by a correct response; if so, we treat the first action as correct. The accuracy of this initial action across different training stages is shown in Fig. 5. As indicated by the dashed green line, accuracy is low in the early stages, reflecting the model’s initial acquisition of search capability, but rises markedly over time. By step 140, roughly 60% of incorrect responses begin with a correct search query, highlighting the high structural similarity shared by first-action actions.
@@
Faster likelihood decay of the first-action response. As illustrated in Fig. 5, we further observe that the likelihood of the first action (light blue) declines substantially more than that of the second action (blue). Early in training, the first-action likelihood is higher, largely because the second action must condition on out-of-distribution feedback. However, as training progresses, both actions exhibit likelihood decay, with the first action eventually dropping below the second around step 110, where the first-action correctness rate is roughly 50%, indicating strong similarity between correct and incorrect trajectories. After step 120, this decay accelerates sharply, marking the onset of the LLD death spiral, where low-likelihood incorrect partial signals further amplify degradation.
Nonsensical first-action. As LLD intensifies and the likelihood of the first action collapses, the model begins producing nonsensical outputs. As illustrated in Fig.˜15 in the appendix, it eventually generates random, meaningless tokens, an effect driven by the severely reduced response likelihood, which causes sampling to select arbitrary words. Although full collapse may not yet have occurred at this stage, the model is already effectively unusable, as it can no longer produce meaningful responses.
As a result, it becomes essential to reduce the unintended penalization of correct actions within incorrect responses. A more detailed discussion is provided in Sec.˜7.
5.4 LLD regularization
To address Lazy Likelihood Displacement (LLD), we introduce a family of likelihood-preserving regularizers that prevent the model from unintentionally reducing the likelihood of responses during GRPO training. Specifically, for the given preserving responses $\bm{y}_{i}\in\mathcal{Y}_{\rm pre}$, we compare its token-level likelihood under the old policy (previous step) and the fine-tuned policy. Our base regularizer penalizes only the tokens whose likelihood decreases after an update.
LLD: Token-Level Likelihood Preservation. LLD applies a token-level penalty: within each response $\mathbf{y}_{i}$, only likelihood-reducing tokens contribute to the loss:
$L_{\rm LLD}=\frac{1}{\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}|\mathbf{y}_{i}|}\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}\sum_{y_{i}\in\mathbf{y}_{i}}\underbrace{\max\left(0,\;\ln\pi_{\theta_{\rm old}}(y_{i}|\mathbf{x},\mathbf{y}_{<i})-\ln\pi_{\theta}(y_{i}|\mathbf{y}_{<i})\right)}_{\text{Likelihood-reducing tokens}}$ (4)
LLDS: Response-Level Gating. To avoid unnecessary penalties on globally improving responses, LLDS introduces a response-level gating mechanism: the penalty activates only when the total likelihood of the response decreases. The LLDS loss is:
$L_{\rm LLDS}=\frac{1}{\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}|\mathbf{y}_{i}|}\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}$ $\underbrace{\mathbbm{1}\!\left[\sum_{y_{i}\in\mathbf{y}_{i}}\left(\ln\pi_{\theta_{\rm old}}(y_{i}|\mathbf{x},\mathbf{y}_{<i})-\ln\pi_{\theta}(y_{i}|\mathbf{y}_{<i})\right)>0\right]}_{\text{Activated only when sum > 0}}\cdot$
$\sum_{y_{i}\in\mathbf{y}_{i}}\underbrace{\max\left(0,\;\ln\pi_{\theta_{\rm old}}(y_{i}|\mathbf{x},\mathbf{y}_{<i})-\ln\pi_{\theta}(y_{i}|\mathbf{y}_{<i})\right)}_{\text{Likelihood-reducing tokens}}$ (5)
This structure preserves normal GRPO learning while directly suppressing responses that suffer from LLD.
LLDS-MA: Masking Answer Tokens. To further encourage multi-step reasoning and tool usage, we mask out final answer tokens from the regularization term. LLDS-MA penalizes likelihood reductions only on reasoning and tool-interaction tokens, excluding the answer span $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{y}_{i,\rm Ans}$}:
$\underbrace{\sum_{y_{i}\in\mathbf{y}_{i}/\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{y}_{i,\rm Ans}}}\max\left(0,\;\ln\pi_{\theta_{\rm old}}(y_{i}|\mathbf{x},\mathbf{y}_{<i})-\ln\pi_{\theta}(y_{i}|\mathbf{y}_{<i})\right)}_{\text{Mask Answer Likelihood-reducing tokens}}$ (6)
Finally, We integrate the regularization into the GRPO objective as
$L_{\rm total}=L_{\rm GRPO}+\lambda L_{\rm{LLDS(-MA)}}$ (7)
where $\lambda$ is the regularization weight. We use LLDS as the default variant and switch to LLDS-MA when stronger encouragement of tool usage is desired. For the preserving set $\mathbf{y}]_{\rm pre}$, we include all responses with non-negative advantages $\hat{A}\geq 0$, ensuring that correct responses ($\hat{A}>0$) and untrained responses ($\hat{A}=0$) do not suffer likelihood reduction. The effect of $\lambda$ is examined empirically in Fig.˜8.
6 Experiments and Analysis
We evaluate the empirical effectiveness of $L_{\rm LLDS(-MA)}$ through comprehensive experiments.
Experimental settings. For training, we follow the setup in jin2025search and conduct experiments using two model families: Qwen-2.5-3B and Qwen-2.5-7B, each in both Base and Instruct variants (yang2024qwen2). We consider two training configurations. (1) NQ-Only (single-hop): the model is trained solely on the single-hop Natural Questions (NQ) dataset (kwiatkowski2019natural). (2) NQ+Hotpot (single-hop+multi-hop): the model is trained on a merged corpus combining NQ with HotpotQA (yang2018hotpotqa), providing broader coverage of both open-domain and multi-hop reasoning. For the retrieval dataset, we use the 2018 Wikipedia dump (karpukhin2020dense) as the knowledge corpus and employ E5 (wang2022text) as the dense retriever. To ensure fair comparison across retrieval-augmented baselines, we fix the number of retrieved passages to three, following the configuration in jin2025search. Unless otherwise specified, we use the same optimization hyperparameters as Search-R1 (jin2025search), with the only modification being a reduced maximum turn limit: from 4 in Search-R1 to 2 for NQ-only training and 3 for the NQ+Hotpot setting, improving training efficiency. Unless stated otherwise, the regularization weight is fixed at $\lambda=0.1$.
Evaluation settings. The Evaluation is performed on the validation or test sets of seven datasets, categorized as follows: (1) General Question Answering: NQ (kwiatkowski2019natural), TriviaQA (joshi2017triviaqa), and PopQA (mallen2022not). (2) Multi-Hop Question Answering: HotpotQA (yang2018hotpotqa), 2WikiMultiHopQA (ho2020constructing), Musique (trivedi2022musique), and Bamboogle (press2022measuring). We adopt exact match (EM) as the primary evaluation metric, consistent with the protocol in jin2025search. Such diversity ensures a thorough examination of search engine-integrated LLM performance across a range of realistic reasoning scenarios.
6.1 Experimental Results
@@
NQ-Only
+LLDS 0.490 0.610 0.450 0.517 0.294 0.193 0.074 0.121 0.171 0.319
+LLDS-MA 0.464 0.616 0.458 0.513 0.363 0.320 0.131 0.315 0.282 0.381 (+13.4%)
NQ+hotpot
+LLDS 0.462 0.622 0.460 0.515 0.432 0.383 0.180 0.395 0.347 0.419 (+24.7%)
Table 1: Results on General QA and Multi-Hop QA datasets for Qwen2.5-3b-Base/Instruct. denotes results from (jin2025search) using NQ+hotpot training. denotes open-domain datasets and denotes retrieval-augmented datasets.
We evaluate our proposed LLD-mitigation strategy on seven open-domain and multi-hop QA benchmarks using two model families: Qwen2.5-3B (Base/Instruct) and Qwen2.5-7B (Base/Instruct). Tables 1 and 2 summarize the EM performance across all settings. Since vanilla GRPO training often collapses and drives the reward to zero, we use the results from Search-R1 jin2025search, corresponding to the best checkpoint before collapse, as the GRPO baseline.
Results on Qwen2.5-3B. As shown in Tab.˜1, GRPO attains an average score of 0.303 in the NQ-only setting. Incorporating LLD raises performance to 0.321 (+5.9%), and LLDS provides a further improvement to 0.323 (+6.6%). In the more challenging NQ+Hotpot setting, LLDS increases the vanilla GRPO score from 0.312 to 0.360, a 15.4% relative gain. Notably, since LLDS tends to limit the base model to invoking search only once, we apply the multi-aspect variant LLDS-MA, which delivers the strongest performance with an average score of 0.430, representing a substantial +37.8% improvement ratio over vanilla GRPO. A similar trend is observed for Qwen2.5-3B-Instruct: LLDS-MA improves performance to 0.381 under NQ-only, and LLDS achieves 0.419 in the NQ+Hotpot setting (+24.7%).
@@
Search-R1 (GRPO) 0.395 0.560 0.388 0.448 0.326 0.297 0.125 0.360 0.277 0.350
NQ-Only
+LLDS 0.514 0.666 0.470 0.550 0.403 0.368 0.151 0.347 0.317 0.417 (+19.1%)
NQ+hotpot
+LLDS 0.504 0.666 0.471 0.547 0.465 0.475 0.209 0.444 0.398 0.462 (+32.0%)
@@
Search-R1 (GRPO) 0.429 0.623 0.427 0.493 0.386 0.346 0.162 0.400 0.324 0.396
NQ-Only
+LLDS 0.517 0.645 0.460 0.541 0.401 0.298 0.158 0.411 0.317 0.413 (+4.3%)
NQ+hotpot
+LLDS 0.498 0.662 0.464 0.541 0.473 0.443 0.228 0.516 0.428 0.469 (+18.4%)
@@
Results on Qwen2.5-7B As shown in Tab.˜2, for Qwen2.5-7B-Base, applying LLDS yields a substantial improvement, reaching an average score of 0.417 when trained on NQ only, representing a 19.1% relative gain over the GRPO baseline trained on NQ+Hotpot. When LLDS is applied under the NQ+Hotpot training setting, performance increases from 0.350 to 0.462, corresponding to an even larger 32.0% improvement. LLDS also achieves the strongest results across nearly all individual datasets, indicating enhanced evidence retrieval and multi-step reasoning capability. For Qwen2.5-7B-Instruct, we observe a similar pattern. Under the NQ+Hotpot training setting, LLDS raises the average score from 0.396 to 0.469 (improved 18.4%). Notably, LLDS achieves the best performance on every Multi-Hop QA benchmark, including 0.473 on 2Wiki, 0.443 on Musique, and 0.516 on Bamboogle.
These results highlight LLDS(-MA) as an effective method for stabilizing GRPO training and substantially enhancing multi-turn reasoning.

### v2

@@ 5.1 Likelihood Dynamic
To characterize the prevalence and evolution of LLD during training, we visualize the training trajectory in Figure˜2. The dynamics exhibit three distinct phases. Phase I (early stagnation). The likelihood of correct responses remains nearly constant despite increasing rewards, indicating the onset of LLD. The subsequent phases correspond to the $\epsilon\leq 0$ regime, where likelihood strictly decreases. Phase II (steady decay). Likelihood declines slowly but consistently, while rewards continue to rise and gradient norms remain stable (e.g., steps 60–120 in Figure˜2). Phase III (acceleration). After a turning point (around step 120 in Figure˜2), likelihood drops sharply and coincides with a rapid increase in gradient magnitude (marked by the red star), leading to gradient explosion and eventual training collapse. We further provide a zoomed-in view showing that, although collapse is ultimately triggered by gradient explosion, instability and reward degradation emerge earlier.
5.2 LLD Death Spiral
In Figure˜2, we observe an accelerated decline in response likelihood as training progresses. We formalize the resulting accelerated likelihood decay as an LLD death spiral.
Definition 5.1 (Informal LLD Death Spiral).
@@ 5.2 LLD Death Spiral
Consider a policy update from $\pi_{\theta_{s}}$ to $\pi_{\theta_{s+1}}$. We say the system enters an LLD Death Spiral if the policy evolution follows the self-reinforcing pattern
$\mathrm{LLD}_{s}\;\Longrightarrow\;C_{s}^{\mathrm{low}}\;\Longrightarrow\;\mathrm{LLD}_{s+1},\qquad\varepsilon_{s+1}<\varepsilon_{s}\leq 0,$
where $\mathrm{LLD}_{s}$ denotes lazy likelihood displacement at iteration $s$, and $C_{s}^{\mathrm{low}}$ denotes low-confidence trajectories whose likelihood decreases under $\pi_{\theta_{s}}$.
A formal definition is provided in Section˜A.2, and the mechanism is visualized in Figure˜3. Beyond the acceleration phase shown in Figure˜2, we further present additional empirical evidence of the LLD death spiral in the following section.
5.3 Empirical Evidence for LLD Death Spiral
@@
We demonstrate the LLD Death Spiral through accelerated entropy explosion and accelerated per-sample LLD.
@@
Accelerated Per-Sample LLD. We conduct a controlled study on the NQ dataset (Kwiatkowski et al., 2019) to examine how GRPO training affect the likelihood of correct responses. Using Qwen2.5-3B-Ins (Yang et al., 2024), we generate eight rollouts per question and retain only samples containing both correct and incorrect responses. For each question, we reinitialize model parameters, apply a single GRPO update, and measure the average log-likelihood change of correct responses:
$\Delta(\bm{x}):=\frac{1}{N^{+}}\sum_{i=1}^{N^{+}}\sum_{t=0}^{T_{i}}\left[\ln\pi_{\theta^{\prime}}(\bm{y}_{i,t}^{+}\mid\bm{x})-\ln\pi_{\theta}(\bm{y}_{i,t}^{+}\mid\bm{x})\right],$ (3)
where $N^{+}$ denotes the number of correct responses. As shown in Figure˜3(a), likelihood degradation emerges as training progresses (orange curve). During the acceleration phase (iterations 120–140), this effect intensifies into an LLD death spiral, with >> 50% samples exhibiting likelihood drops.
Figure 4: We visualize the evolution of entropy, response length, and valid-search ratio during training. For both Qwen2.5-3B-Instruct and Qwen2.5-3B-Base, entropy rises sharply prior to collapse, indicating severe likelihood displacement.
@@
Accelerated Entropy Explosion. Figure˜4 shows the evolution of entropy, response length, and valid-search ratio for Qwen2.5-3B-Ins and Qwen2.5-3B-Base. For both models, token entropy increases slowly in early training, consistent with the slow-LLD regime in Figure˜2, and then accelerates sharply during the LLD acceleration phase. This entropy surge reflects the LLD death spiral: accumulating low-confidence responses induce increasingly diffuse token distributions, further amplifying likelihood decay and instability. Notably, response length and valid-search frequency remain nearly constant, indicating that the entropy growth and likelihood degradation are driven by LLD rather than longer trajectories or increased tool usage.
5.4 Why Tool-Integrated (TI) GRPO Amplifies LLD
@@
We then show that the severity of LLD in TI-GRPO stems from two interacting factors predicted by Theorem˜4.2 and demonstrated in Figure˜3(b): high response similarity between correct and incorrect trajectories, and low likelihood responses that induce large prediction-error weights.
Frequent correct actions in incorrect responses. Our analysis on Qwen2.5-3B-Ins that trained on the NQ dataset reveals a striking pattern: correct actions frequently appear within otherwise incorrect responses. Specifically, in TI GRPO, the first response action typically formulates a search query, whose correctness increases steadily during training regardless of final answer. We label the first action of incorrect responses as correct if its retrieved documents match those of a correct response. As shown in Fig. 5, this accuracy (green line) is initially low but rises sharply, reaching $\sim$60% by step 140, indicating high similarity on the first action between correct and incorrect trajectories. At the same time, the likelihood of the first action (light blue) decays faster than that of the second action (blue): although initially higher due to the second action conditioning on OOD feedback, it drops below the second action around step 110, where first-action correctness reaches $\sim$50%. After step 120, this decay accelerates sharply, where low-likelihood incorrect prefixes further amplify degradation. As LLD intensifies, the likelihood of the first action collapses, and the model begins producing random, nonsensical tokens (Appendix, Fig. 15), rendering responses effectively unusable even before full training collapse. These results motivate mitigating unintended penalization of correct actions within incorrect trajectories (see Appendix˜D).
Figure 6: Effect of off-policy tool context on negative log-likelihood. Increasing the number of tool-interaction rounds consistently reduces response likelihood.
Low Likelihood reasoning after tool feedback. A second factor that amplifies LLD is the use of OOD tool feedback as a conditioning context. They introduce a persistent distribution mismatch between the policy and its input context, systematically reducing the likelihood assigned to subsequent actions and increasing negative log-likelihood. As shown in Figure˜6, computed on the first 50 NQ training samples, this effect strengthens with the number of tool-interaction rounds, producing a monotonic increase in NLL for both Qwen2.5-3B-Instruct and Qwen2.5-7B-Instruct models.
5.5 LLD Suppression regularization
To mitigate LLD, we propose a likelihood-preserving regularizer, $\rm LLDS$, which prevents unintended reductions in response likelihood during GRPO training. Specifically, for each given preserving response $\bm{y}_{i}\in\mathcal{Y}_{\rm pre}$, we compare token-level likelihoods under the previous and current policies, and quantify their change by
$\Delta_{i,t,k}$ $=\ln\pi_{\theta_{\mathrm{old}}}\!\left(y_{i,t,k}\mid\mathbf{x},\bm{y}_{i,<t},\bm{o}_{i,<t},\bm{y}_{i,t,<k}\right)$
$-\ln\pi_{\theta}\!\left(y_{i,t,k}\mid\mathbf{x},\bm{y}_{i,<t},\bm{o}_{i,<t},\bm{y}_{i,t,<k}\right).$
where $i$ indexes samples, $t$ actions (turns), and $k$ tokens. Positive $\Delta_{i,t,k}$ indicates a likelihood decrease after the update. These tokens are candidates for regularization.
@@ 5.5 LLD Suppression regularization
ZeroSearch-Base 0.430 0.616 0.414 0.487 0.338 0.346 0.130 0.139 0.238 0.345
ZeroSearch-Ins 0.414 0.574 0.448 0.479 0.274 0.300 0.098 0.111 0.196 0.317
StepSearch 0.296 0.490 0.341 0.375 0.294 0.360 0.140 0.258 0.263 0.311
CriticSearch - - - - 0.414 0.409 0.180 0.368 0.343 -
Tree-GRPO 0.468 0.597 0.436 0.500 0.424 0.437 0.178 0.432 0.368 0.424
@@
+LLDS 0.462 0.627 0.468 0.519 0.443 0.447 0.197 0.444 0.383 0.441 (+31.3%)
Search-R1-GSPO 0.376 0.561 0.374 0.437 0.332 0.329 0.098 0.250 0.252 0.331
+LLDS 0.489 0.640 0.483 0.537 0.466 0.456 0.215 0.403 0.385 0.451(+36.3%)
Table 1: Results on General QA and Multi-Hop QA datasets for Qwen2.5-3b-Base/Instruct. All results are reported using Exact Match (EM). denotes results from (Jin et al., 2025). denotes results from (Zhang et al., 2025) that only have multi-hop QA results. denotes in-distribution datasets, while denotes out-of-distribution datasets.
$\rm LLDS$: Action-Level Gating. To avoid unnecessary penalties on globally improving responses, we introduce an action-level gating mechanism: the penalty activates only when the total likelihood of an action decreases. The resulting $\rm LLDS$ loss is defined as:
$L_{\mathrm{LLDS}}$ $=\frac{1}{\sum_{i=1}^{|\mathcal{Y}_{\mathrm{pre}}|}\sum_{t=0}^{T_{i}}|\bm{y}_{i,t}|}\sum_{i=1}^{|\mathcal{Y}_{\mathrm{pre}}|}\sum_{t=0}^{T_{i}}$
$\underbrace{\mathbbm{1}\!\left[\sum_{k=1}^{|\bm{y}_{i,t}|}\Delta_{i,t,k}>0\right]}_{\text{Activated only when sum > 0}}\sum_{k=1}^{|\bm{y}_{i,t}|}\underbrace{\max(0,\Delta_{i,t,k})}_{\text{Likelihood-reducing tokens}},$ (4)
This structure directly suppresses actions that suffer from LLD. We also introduce token-level and response-level variants in Section˜B.3, unless stated otherwise, $\rm LLDS$ refers to the action-level gating.
LLDS-MA: Masking Answer Tokens. To further promote multi-turn reasoning and tool usage, we assign smaller regularization weights to final answer tokens, discouraging premature answer generation. Consequently, $\rm LLDS$-MA penalizes likelihood reductions on reasoning and tool-interaction tokens more strongly, while only softly regularizing the answer span $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{y}_{i,\rm Ans}$}:
$\sum_{k=1}^{|\mathbf{y}_{i,t}|}w_{i,t,k}\;\max\!\left(0,\;\Delta_{i,t,k}\right),w_{i,t,k}=\begin{cases}\beta,&k\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{y}_{i,\rm Ans}}\\ 1,&\text{otherwise}\end{cases}$ (5)
Finally, we integrate the regularization into the GRPO objective as
$L_{\rm total}=L_{\rm GRPO}+\lambda L_{\rm{LLDS(-MA)}}$ (6)
where $\lambda$ is the regularization weight. We use LLDS as the default variant, and adopt LLDS-MA when the model collapses to single-turn behavior. The preserving set $\mathcal{Y}_{\rm pre}$ includes all responses with non-negative advantages $\hat{A}\geq 0$, ensuring that correct responses ($\hat{A}>0$) and untrained responses ($\hat{A}=0$) do not suffer LLD. The effect of $\lambda$ is examined empirically in Figure˜8.
@@
ZeroSearch-Base 0.424 0.664 0.604 0.564 0.320 0.340 0.180 0.333 0.293 0.409
ZeroSearch-Ins 0.436 0.652 0.488 0.525 0.346 0.352 0.184 0.278 0.290 0.391
StepSearch 0.364 0.565 0.390 0.440 0.346 0.394 0.183 0.444 0.342 0.382
CriticSearch - - - - 0.442 0.428 0.194 0.472 0.384 -
Tree-GRPO 0.481 0.633 0.452 0.522 0.446 0.423 0.202 0.440 0.378 0.439
@@
Search-R1-GRPO 0.395 0.560 0.388 0.448 0.326 0.297 0.125 0.360 0.277 0.350
+LLDS 0.512 0.672 0.480 0.555 0.486 0.469 0.232 0.508 0.424 0.480 (+37.1%)
@@
Search-R1-GRPO 0.429 0.623 0.427 0.493 0.386 0.346 0.162 0.400 0.324 0.396
+LLDS 0.518 0.676 0.486 0.560 0.492 0.490 0.228 0.492 0.426 0.483 (+22.0%)
@@
6 Experiments and Analysis
Experimental settings. For training, we follow the setup in (Jin et al., 2025) and conduct experiments using two model families: Qwen-2.5-3B and Qwen-2.5-7B, each in both Base and Instruct variants (Yang et al., 2024). We follow the training configuration in Search-R1 (Jin et al., 2025) and use NQ+Hotpot (single-hop+multi-hop): the model is trained on a merged corpus combining NQ with HotpotQA (Yang et al., 2018), providing broader coverage of both open-domain and multi-hop reasoning. For the retrieval dataset, we use the 2018 Wikipedia dump (Karpukhin et al., 2020) as the knowledge corpus and employ E5 (Wang et al., 2022) as the dense retriever and fix the number of retrieved passages to three. Unless otherwise specified, we use the same optimization hyperparameters as Search-R1 (Jin et al., 2025), with the only modification being a reduced maximum turn from 4 to 3 for the NQ+Hotpot, improving training efficiency. Unless stated otherwise, the regularization weight is fixed at $\lambda=0.2$. More details see Sec. B.2.1.
Evaluation settings. The Evaluation is performed on seven datasets, categorized as follows: (1) General QA: NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and PopQA (Mallen et al., 2022). (2) Multi-Hop QA: HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020), Musique (Trivedi et al., 2022), and Bamboogle (Press et al., 2022). We follow (Jin et al., 2025) and adopt exact match (EM) as the evaluation metric.
Baseline Methods. To evaluate the effectiveness of $\rm LLDS$, we use baselines in Search-R1 (Jin et al., 2025) and additionally include methods that improve training stability via dense rewards, namely StepSearch (Zheng et al., 2025b), CriticSearch (Zhang et al., 2025), and TreeGRPO (Ji et al., 2025). See Section˜B.2.2 for detailed descriptions.
Figure 7: Comparisons of training reward curves between baseline GRPO (blue) and GRPO + LLDS (red). From left to right: Qwen-2.5-3B-Base, Qwen-2.5-3B-Instruct, Qwen-2.5-7B-Base, and Qwen-2.5-7B-Instruct. The baseline consistently collapses (reward drops to zero), while LLDS stabilizes training and sustains high rewards.
6.1 Experimental Results
We evaluate our proposed $\rm LLDS$ on seven open-domain and multi-hop QA benchmarks using two model families: Qwen2.5-3B (Base/Instruct) and Qwen2.5-7B (Base/Instruct). Tables 1 and 2 summarize the EM performance across all settings. Since vanilla GRPO training often collapses, we use the results from Search-R1 (Jin et al., 2025), corresponding to the best checkpoint before collapse, as the GRPO baseline. This choice favors GRPO and avoids underestimating its performance. More results see Section˜B.4.
Results on Qwen2.5-3B. As shown in Table˜1, $\rm LLDS$ improves the vanilla GRPO score from 0.312 to 0.360, corresponding to a 15.4% relative gain. Notably, because the 3B base model tends to invoke search only once, we apply $\rm LLDS$-MA, which achieves the best performance with an average score of 0.440, representing a substantial +45.2% improvement over GRPO. A similar trend is observed for Qwen2.5-3B-Instruct, where LLDS attains 0.441 under the NQ+Hotpot setting (+31.3%).
Results on Qwen2.5-7B. As shown in Table˜2, $\rm LLDS$ delivers substantial gains over vanilla GRPO for both base and instruct variants. For Qwen2.5-7B-Base, it improves the average score from 0.350 to 0.480, a 37.1% relative increase, while for Qwen2.5-7B-Instruct, performance rises from 0.396 to 0.483, corresponding to a 22.0% gain. Notably, $\rm LLDS$ attains the best results on all Multi-Hop QA benchmarks with MH-Avg accuracy being 0.426.
Comparison with Dense Rewards. As shown in Tables˜1 and 2, stabilizing outcome-reward GRPO training with $\rm LLDS$ consistently outperforms methods that rely on dense reward supervision. These dense-reward approaches typically require turn-level ground-truth signals, external LLM judges, or more expensive tree-based rollouts. In contrast, $\rm LLDS$ achieves superior performance using only outcome rewards. Notably, $\rm LLDS$ outperforms the strongest dense-reward baseline, Tree-GRPO, by 4% and 10% on the 3B and 7B models, respectively, and exceeds other dense-reward methods by over 30%, demonstrating both improved effectiveness and substantially lower supervision.
[…]